OpenStudy (anonymous):

Find a rational number and an irrational number that are between 9.5 and 9.7.

OpenStudy (igreen):

Rational numbers are numbers that can be written as fractions. 9.6 can be written as \(9 \dfrac{3}{5}\) or \(\dfrac{48}{5}\). So 9.6 is your rational number.

OpenStudy (igreen):

Irrational numbers cannot be written as fractions. One example is \(\pi\). It goes on forever but does NOT repeat. For your irrational number you can have 9.562359123571283578... etc. It goes on forever but does NOT repeat.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Kittymeowb

OpenStudy (anonymous):

Thank you! cx

OpenStudy (igreen):

Np.